Tan 30 degree - Cot 30 degree
Answer by drk(1908)

Tan(3) =1/sqrt(3) = sqrt(3)/3
cot(3) = sqrt(3)
--
sqrt(3)/3 - srt(3) = -2*sqrt(3)/3